Private colleges and universities rely on money contributed by individuals and corporations for their operating expenses. Much of this money is put into a fund called an endowment, and the college spends only the interest earned by the fund. A recent survey of 8 private colleges in the United States revealed the following endowments (in millions of dollars): 60.2, 47.0, 235.1, 490.0, 122.6, 177.5, 95.4, and 220.0. Summary statistics yield x = 180.975 and s = 143.042. A 95% confidence interval for the mean endowment of all the private colleges in the United States is:
Answer #1

Given that, sample size (n) = 8, sample mean = 180.975 and

sample standard deviation (s) = 143.042

Degrees of freedom = 8 - 1 = 7

t-critical value at significance level of 0.05 with 7 degrees of freedom is,

Excel Command : =TINV(0.05, 7) = 2.365

The 95% confidence interval for the population mean is,

Therefore, a 95% confidence interval for the mean endowment of all the private colleges in the United States is (61.3699, 300.5801).